OpenStudy (anonymous):

Space junk falls from rest toward the Earth from a height of 690 km above the Earth’s surface. G= 6.673E-11 N.m ²/kg ²; M_(earth)= 5.98E24kg; R_(earth) = 6,380,000m; You may assume that the junk encounters almost no air resistance above a height of approximately 20 km. What is the speed of the junk as it falls to a height of 32.8 km?

OpenStudy (anonymous):

any idea as how you may proceed with this problem?!

OpenStudy (anonymous):

so confusing so many heights

OpenStudy (anonymous):

Not really.. when u do gravitation.. u always deal with such data.. :)

OpenStudy (anonymous):

any formula?

OpenStudy (anonymous):

\[v =\sqrt{\frac{ GM }{ R }}\]

OpenStudy (anonymous):

M I RIGHT?

OpenStudy (anonymous):

I was thinking that also, but isn't that for the speed required for a satellite to orbit the Earth without falling?

OpenStudy (anonymous):

\[v =\sqrt{g[h1-h2]}\]

OpenStudy (anonymous):

@Mashy where r u?we need help

OpenStudy (anonymous):

vfinal2 = vinitial2 + [2g (sfinal - sinitial)] The square of the final velocity equals the square of the initial velocity plus the product of twice the acceleration due to gravity times the difference in the altitudes or heights of the falling body from beginning to end.

OpenStudy (anonymous):

Whats 'g' in that equation? Gravitational acceleration? doesn't this change with height?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

How would I go about using that equation to solve for velocity when acceleration is always changing :s

OpenStudy (anonymous):

Hang on, I think I've found some relevant information. I think I will have to use work and energy procedures. I'll post back if I'm stuck or I get the question >.<

OpenStudy (anonymous):

Oh nevermind no given mass.

OpenStudy (anonymous):

What would i substitute in g using vfinal2 = vinitial2 + [2g (sfinal - sinitial)]

OpenStudy (jt950):